# Towards A Universal Graph Structural Encoder

### v1

@@ 3.4 Expressive Power of GFSE
We show that GFSE can generate highly expressive PSE by incorporating relative edge encoding into the attention computation in the Graph Transformer backbone. Specifically, we employ the Structural Encoding enhanced Global Weisfeiler-Lehman test (SEG-WL) (Zhu et al., 2023), a generalized WL test that incorporates relative structural encoding into the isomorphism algorithm, to characterize the expressiveness of GFSE. For an input graph $G(V,E)$ with node set $V$ and edge set $E$, let $f_{P}:V\rightarrow\mathcal{X}$ and $f_{R}:V\times V\rightarrow\mathcal{X}$ indicate the node-level and edge-level structural encoding, respectively. Different from traditional WL test, SEG-WL updates the node labels at the $t$-th iteration by $g_{t}(v)=\text{hash}\left(\left\{\left\{\left(g_{t-1}(u),f_{R}(v,u)\right):u\in V\right\}\right\}\right)$ and $g_{0}(v)=\text{hash}(f_{P}(v))$. SEG-WL can be viewed as a high-level abstraction of the learning paradigm of our pre-training architecture with biased attention (Eq. 3), where relative structural encoding between any two nodes is considered for updating node representations (Zhu et al., 2023). Let RW($d$)-SEG-WL denote the case that $f_{P}$ and $f_{R}$ are determined by $\mathbf{P}$ and $\mathbf{R}$ with $d$ dimension, i.e., $f_{P}(v_{i})=\mathbf{P}_{i}\in\mathbb{R}^{d}$ and $f_{R}(v_{i},v_{j})=\mathbf{R}_{ij}\in\mathbb{R}^{d}$ for the $i$-th and $j$-th nodes. We have the following propositions.
@@
The theoretical proof and empirical verification are given in Appendix D. RW-SEG-WL is able to distinguish all low-order graphs with orders equal to or less than 8 and successfully distinguishes most strongly regular graphs where 3-WL fails to distinguish. RW-SEG-WL stands as an expressivity upper bound of our proposed GFSE. The pre-training tasks are meticulously designed to push GFSE towards achieving the upper bound established by RW-SEG-WL. These pre-training tasks optimize both node-level and edge-level structural encoding, progressively refining the effectiveness of the model in generating expressive encoding.
3.5 Computational Complexity
GFSE’s complexity comprises two parts: pre-computation of self-supervision labels and pre-training. Computing the shortest path distance between node pairs results in the time complexity of $\mathcal{O}(|E|+|V|\log|V|)$ with node set $V$ and edge set $E$. A brute-force implementation of the subgraph isomorphism counting of fixed size $t$ is $\mathcal{O}(|V|^{t})$. We consider the graphlets with at most 5 nodes. One can also choose special graphlet types, e.g., paths, cycles, and triangles, which can be efficiently enumerated (Giscard et al., 2019). Approximating and scalable algorithms can be further used to accelerate this pre-processing step (Fu et al., 2024; Ying et al., 2020; Pashanasangi & Seshadhri, 2020). For pre-training, the complexity is $\mathcal{O}(|V|^{2})$ for full attention computation and $\mathcal{O}(d|V||E|)$ for initial encoding computation. See runtime evaluations in Appendix F.2.
4 Experiments
We first evaluate the pre-training performance in Sec. 4.2 and empirically assess the expressiveness of GFSE on synthetic datasets (Sec. 4.3). We then evaluate GFSE in a wide range of downstream graph learning tasks in Sec. 4.4. Specifically, we conduct experiments with LLMs on text-attributed graphs in Sec. 4.5.
4.1 Pre-Training Setup
Dataset. We utilize diverse datasets for pretraining, ensuring a broad spectrum of graph structures and scales, including MolPCBA, MolHIV, MNIST, peptides, ogbn-proteins, Pokec, ogbn-arxiv and ogbn-product (Wu et al., 2018; Bhatia et al., 2016; Mikolov et al., 2013; Szklarczyk et al., 2019; Chiang et al., 2019; Takac & Zabovsky, 2012; Dwivedi et al., 2023, 2022). These datasets cover several real-world graph domains, such as social networks, academic networks, etc. We refer to Table 6 in Appendix C.1 for detailed dataset statistics. For large-scale graphs, we first partition them into sets of subgraphs by the METIS algorithm (Karypis & Kumar, 1997) to handle scalability issues. Training samples from different datasets are mixed and randomly shuffled to form a large-scale pre-training dataset.
Training Setting. The pre-training stage is conducted on the standard train/validation/test splits of the pre-training datasets. The initial encoding dimension $d$ is set as 8. We adopt GIN (Xu et al., 2018) as the message-passing layer in the GPS and adopt 8 GPS layers with 8 heads and 128 hidden dimensions. The output dimension is 64 by default. We use Adam as the optimizer with an initial learning rate of $0.001$ and the batch size is 256. The maximum training epochs is 100. An early stopping strategy is used to mitigate overfitting. The pre-training is implemented on the NVIDIA A40 48GB GPU. We refer to Appendix C.5 for more details. The code is available at https://github.com/Graph-and-Geometric-Learning/GFSE.
4.2 Pre-training Evaluation
GFSE is pre-trained with four self-supervised learning tasks. We iteratively change the message passing layers (e.g., GatedGCN (Bresson & Laurent, 2017), GCN (Kipf & Welling, 2016a) and GIN (Xu et al., 2018)) and replace the biased attention with traditional self-attention in the default GFSE architecture. We evaluate the pre-training performance with different architectures as shown in Figure 2. Accuracy is used to measure community detection and graph contrastive learning tasks, indicating the proportion of node (graph) pairs that are correctly predicted. MSE and MAE are used for shortest path distance and motif counting tasks. See more details in Appendix C.2.
Figure 2: Pre-training performance with different architectures. TF is the abbreviation of transformer.
We observe a consistent performance boost when combining biased attention with GIN, which serves as the default configuration of GFSE. The pre-trained GFSE is able to generate effective PSE, which can reconstruct other pre-defined PSE, such as LapPE (Lim et al., 2022) and ElstaticPE (Kreuzer et al., 2021). We provide experimental results in Appendix D.3. The ability to reconstruct various types of PSE, without taking them as training objectives explicitly, demonstrates the sufficiency and effectiveness of the chosen tasks in pre-training the model for expressive and generalizable graph representations.
4.3 Expressiveness Evaluation
We empirically evaluate the structure-awareness of the positional and structural encoding (PSE) generated by GFSE on Triangle (Knyazev et al., 2019), Pattern, and Cluster (Dwivedi et al., 2023) , three datasets that require discerning intricate graph topologies. We evaluate the performance boost brought by GFSE in comparison to traditional RWSE and LapPE. We test on various downstream graph learning models, including MLP, GIN (Xu et al., 2018), transformer (Vaswani et al., 2017) and GPS (Rampášek et al., 2022). We further compare with the learning-based PSE approach: GPSE (Liu et al., 2023b). See Appendix A for more baseline details. As shown in Table 1, GFSE generates expressive and robust PSE that consistently improves the base model’s performance, whereas other structural encodings exhibit considerable variation across different datasets or different base models. Notably, the performance boost brought by GFSE is particularly significant for the Transformer on the Triangle dataset. This demonstrates that GFSE has a stronger enhancement effect on models that originally lack structural bias.
@@ 4.3 Expressiveness Evaluation
4.4 Downstream Evaluation
4.4.1 Evaluation on Pre-training Dataset
Dataset and Baseline. We conduct a comprehensive evaluation of GFSE on real-world graph datasets. We adhere to the experimental setting and hyper-parameters established by previous works (Rampášek et al., 2022) to implement base models. We augment the initial node features with GFSE and evaluate the performance in downstream tasks, comparing it against established structural encoding methods including RWSE, LapPE, and GPSE (Liu et al., 2023b).
[…]
Table 2: Performance on MolPCBA, ZINC (subset), Peptides-func and Peptides-struct.
MolPCBA ZINC Peptides-func Peptides-struct
AP $\uparrow$ MAE $\downarrow$ AP $\uparrow$ MAE $\downarrow$
GIN+LapPE $0.2701_{\pm 0.0013}$ $0.2203_{\pm 0.0386}$ $0.5323_{\pm 0.0083}$ $0.2650_{\pm 0.0041}$
GPS+RWSE $0.2907_{\pm 0.0028}$ $0.0700_{\pm 0.0040}$ $0.6603_{\pm 0.0085}$ $0.2739_{\pm 0.0063}$
GPS+GPSE $0.2911_{\pm 0.0036}$ $0.0648_{\pm 0.0030}$ $0.6688_{\pm 0.0151}$ $\mathbf{0.2464}_{\pm 0.0025}$
GPS+GFSE $0.2916_{\pm 0.0061}$ $\mathbf{0.0613}_{\pm 0.0026}$ $\mathbf{0.6874}_{\pm 0.0120}$ $0.2474_{\pm 0.0051}$
[…]
4.4.2 Evaluation on Molecular Graphs
Settings and Methods. We use the molecular datasets namely Tox21, Sider, BBBP, and MUV (Hu et al., 2020a) as a downstream task. We evaluate the effectiveness of GFSE under two settings: (1) In the training from scratch setting, we directly concatenate GFSE’s PSE with the raw node features to create new input features. This augmented representation is then fed into a randomly initialized model from the beginning of training. We take GINE (Xu et al., 2018) and GPS (Rampášek et al., 2022) as our backbone. (2) In the fine-tuning setting, we assess GFSE’s ability to enhance pre-trained models by concatenating the node encodings obtained from a pre-trained model with the GFSE’s PSE. The concatenated features are then fed into the final read-out layers for prediction. During fine-tuning, the parameters of the pre-trained model and read-out layers are updated. We select GraphMAE (Hou et al., 2022) and MoleBERT (Xia et al., 2022) as the pre-trained backbones and compare with other baselines without structural encoding, namely SSP (Hu et al., 2019), GraphLoG (Xu et al., 2021b), GraphCL (You et al., 2020). Refer to Appendix E.1 for more implementation details.

### v2

@@ 3.2. Expressive Power of GFSE
We show that GFSE can generate highly expressive PSE by incorporating relative edge encoding into the attention computation. To characterize the expressiveness of GFSE, we employ the Structural Encoding enhanced Global Weisfeiler-Lehman test (SEG-WL) (zhu2023structural), a generalized WL test that incorporates relative structural encoding into the isomorphism algorithm. For an input graph $G(V,E)$ with node set $V$ and edge set $E$, let $f_{P}:V\rightarrow\mathcal{X}$ and $f_{R}:V\times V\rightarrow\mathcal{X}$ indicate the node-level and edge-level structural encoding, respectively. Different from traditional WL test, SEG-WL updates the node labels at the $t$-th iteration by $g_{t}(v)=\text{hash}\left(\left\{\left\{\left(g_{t-1}(u),f_{R}(v,u)\right):u\in V\right\}\right\}\right)$ and $g_{0}(v)=\text{hash}(f_{P}(v))$. SEG-WL can be viewed as a high-level abstraction of the learning paradigm of our pre-training architecture with biased attention (Eq. 2), where relative structural encoding between any two nodes is considered for updating node representations (zhu2023structural). Let RW($d$)-SEG-WL denote the case that $f_{P}$ and $f_{R}$ are determined by $\mathbf{P}$ and $\mathbf{R}$ with $d$ dimension, i.e., $f_{P}(v_{i})=\mathbf{P}_{i}\in\mathbb{R}^{d}$ and $f_{R}(v_{i},v_{j})=\mathbf{R}_{ij}\in\mathbb{R}^{d}$ for the $i$-th and $j$-th nodes. We have the following propositions.
@@
Proof.
We first introduce Neighbor-SEG-WL, which is the SEG-WL test when $f_{P}$ is an identity encoding and $f_{R}(u,v)$ equals 1 if $(u,v)\in E$ and $2$ otherwise. Previous works have proved the following Proposition (zhu2023structural):
Lemma 0.
Therefore, Neighbor-SEG-WL is a specific example of SEG-WL test that has equivalent expressiveness to the 1-WL test. We then prove that RW-SEG-WL is strictly more expressive than Neighbor-SEG-WL. Let $d_{\text{neg}}(u,v)$ indicate the edge-level encoding $f_{R}$ in Neighbor-SEG-WL. Note $d_{\text{neg}}(v_{i},v_{j})=2$ if and only if $A_{ij}=0$. Recall that $f_{R}(\cdot,\cdot)$ in RW-SEG-WL satisfies $f_{R}(v_{i},v_{j})=\mathbf{R}_{ij}\in\mathbb{R}^{d}$ with $\mathbf{R}=[\mathbf{I},\mathbf{M},\cdots,\mathbf{M}^{d-1}]$ where $\mathbf{M}=\mathbf{D}^{-1}\mathbf{A}$. Therefore, $f_{R}$ in RW-SEG-WL strictly contains the information of $d_{\text{neg}}$. Therefore, if two non-isomorphic graphs can be distinguished by WL, they can be distinguished by RW-SEG-WL. Proposition 3.1 is proved. ∎
Figure 2. RW($d$)-SEG-WL can distinguish two non-isomorphic graphs when $d>4$ while 3-WL fails.
@@
Proof.
To prove Proposition 3.3, we provide an example in Figure 2 which shows the Shrikhande graph and the Rook’s $4\times 4$ graph, a pair of strongly regular graphs. It has been proved that they cannot be distinguished by 3-WL (arvind2020weisfeiler). We empirically verified that RW($d$)-SEG-WL with $d>4$ can distinguish these two graphs. ∎
Table 1. Results of synthetic graph isomorphism tests
# graphs 21 112 853 11117 15 10 41 180 28 78
# graph pairs 210 6216 363378 61788286 105 45 820 16110 378 3003
number of undistinguishable graph pairs
WL 0 3 17 312 105 45 820 16110 378 3003
SPD-SEG-WL 0 2 12 186 105 45 820 16110 378 3003
RW-SEG-WL($d=8$) 0 0 0 0 0 0 0 0 0 0
Synthetic Graph Isomorphism Tests. To evaluate the expressive power of RW-SEG-WL, we perform synthetic graph isomorphism tests on low-order graphs and strongly regular graphs. We consider low-order graphs with up to 8 nodes. Strongly regular graph pair SRG($n,k,\lambda,\mu)$ means graphs with $n$ nodes, where each node has $k$ neighbors. Each adjacent pair of nodes has the same number $\lambda$ of neighbors in common, and each non-adjacent node pair has $\mu$ neighbors in common. Strongly regular graphs are known to be challenging cases for graph isomorphism test algorithms due to their highly symmetric structure (arvind2020weisfeiler). We compare with 1-WL and SPD-SEG-WL, where $f_{P}$ is an identity encoding and $f_{R}$ is defined as the shortest path distance between two nodes. Note SPE-SEG-WL can be viewed as an expressivity upper bound of Graphormer (ying2021transformers). The results are shown in Table 1. We observe that RW-SEG-WL can distinguish significantly more non-isomorphic graphs than 1-WL and SPD-SEG-WL. Specifically, with $d$ equals 8, i.e., when considering a random walk with 8 steps, RW-SEG-WL successfully distinguishes all low-order graphs and strongly regular graphs. When setting $d=4$, there are 16 pairs of strongly regular graphs that cannot be distinguished by RW-SEG-WL. Therefore, it is natural to develop a graph transformer equipped with a relative random-walk encoding that can accurately capture important graph structures and demonstrate strong expressive power.
@@
RW-SEG-WL stands as an expressivity upper bound of our proposed GFSE. The pre-training tasks are meticulously designed to push GFSE towards achieving the upper bound established by RW-SEG-WL. The pre-training tasks are carefully designed to optimize both node-level and edge-level structural encoding, progressively refining the effectiveness of the model in generating expressive encoding.
3.3. Self-supervised Pre-training Tasks
GFSE is pre-trained with reconstruction and contrastive structural tasks at multiple context levels. Each task highlights a specific structural aspect, thereby augmenting the model’s expressiveness and capability to capture complex graph structures. Let $\mathbf{P}^{L}\in\mathbb{R}^{N\times d_{e}}$ represent the output after $L$ layers. We decode $\mathbf{P}^{L}$ with independent MLP heads for each pre-training task.
Shortest Path Distance Regression encodes the global proximity and connectivity between nodes, which helps to discern nodes’ positions and relations within the entire graph (li2020distance). We pre-compute the shortest path distance via the Dijkstra algorithm (dijkstra1959note) to create the label $\text{SPD}\in\mathbb{R}^{N\times N}$. The loss for shortest path distance regression is computed by $\mathcal{L}_{\text{SPD}}=\frac{1}{|E|}\sum_{i,j\in V}\|h_{\text{SPD}}(\mathbf{P}^{L}_{i}\|\mathbf{P}^{L}_{j})-\text{SPD}_{i,j}\|^{2}$, where $\|$ indicates the concatenation and $h_{\text{SPD}}$ indicates a task-specific head.
Motif Counting enables the model to better identify each node’s role in the surrounding subgraphs. To improve expressiveness, we include a variety of small motifs, called graphlets (prvzulj2004modeling; prvzulj2007biological; bouritsas2022improving), with different numbers of vertices, which are beyond usual types like stars, paths, cycles, and cliques. Let $Y_{i}^{m}\in\mathbb{Z}^{k}$ denote the motif label that indicates the number of certain motifs surrounding node $i$, where $k$ is the number of graphlet types. The loss is formulated as $\mathcal{L}_{\text{MC}}=\frac{1}{|V|}\sum_{i\in V}\|h_{\text{MC}}(\mathbf{P}^{L}_{i})-Y_{i}^{m}\|^{2}$, where $h_{\text{MC}}:\mathbb{R}^{d}_{e}\rightarrow\mathbb{R}^{k}$ is the task-specific head for motif counting.
Community Detection aims to identify densely connected subgraphs, where nodes within a community are more closely linked to each other than to nodes outside the community. Such community structures are ubiquitous in various real-world networks, e.g., social networks, and transportation systems. We employ the Louvain Community Detection Algorithm (blondel2008fast) to extract the community structure from pre-training graphs, which clusters nodes into communities based solely on graph topology without node features. We approach this task in a contrastive learning manner by minimizing the embedding distances between intra-community nodes while maximizing the distance between inter-community nodes.
$\mathcal{L}_{\text{CD}}=\sum_{i\in V}\sum_{j\in V}Y_{i,j}^{c}(1-\operatorname{sim}(i,j))+(1-Y_{i,j}^{c})\max(0,\epsilon-(1-\operatorname{sim}(i,j))),\vskip-5.69046pt$
@@ 3.3. Self-supervised Pre-training Tasks
Graph Contrastive Learning aims to distinguish graphs from different datasets. The motivation is rooted in the observation that similar structural characteristics in different domains may exhibit distinct meanings. For example, a densely connected subgraph in a citation network might correspond to a specific research topic, whereas a similar structure in a social network could represent a tightly-knit community or interest group. Therefore, GFSE distinguishes graphs from different datasets with the contrastive loss $\mathcal{L}_{\text{GCL}}=-\log\frac{\exp\left(\operatorname{sim}\left(\boldsymbol{z}_{G_{i}},\boldsymbol{z}_{G_{j}}\right)/\tau\right)}{\sum_{k=1}^{K}\mathbbm{1}_{[G_{k}\nsim G_{i}]}\exp\left(\operatorname{sim}\left(\boldsymbol{z}_{G_{i}},\boldsymbol{z}_{G_{k}}\right)/\tau\right)}$, where $\tau$ is the temperature, $G_{i}$ and $G_{j}$ are from the same dataset, $\boldsymbol{z}_{G_{i}}=\mathrm{GlobalPool}(h_{\text{GCL}}(\mathbf{P}^{L}_{G_{i}}))$ is the output of the global pooling applied to the last layer’s representation $\mathbf{P}^{L}_{G_{i}}$ for the graph $G_{i}$, $K$ is the number of negative samples, and $\mathbbm{1}_{[G_{k}\nsim G_{i}]}$ indicates whether graphs $G_{k}$ and $G_{i}$ originate from different datasets.
Multi-task Selection and Loss Balance: Each pre-training task targets a different structural aspect, enabling GFSE to capture a comprehensive understanding of graph topology. For instance, the shortest path distance regression task emphasizes global connectivity within graphs, while motif counting focuses on the occurrence of local subgraph patterns. Although other pre-training tasks could be considered, the chosen tasks provide a balanced and effective combination, capable of capturing both local and global structural information at varying granularities without introducing unnecessary complexity. Since the loss scale and difficulty vary significantly across tasks, we employ task-specific uncertainty (kendall2018multi) to unify the loss scales, which automatically balances different pre-training losses, i.e., $\mathcal{L}_{\text{SPD}},\mathcal{L}_{\text{MC}},\mathcal{L}_{\text{CD}}$, and $\mathcal{L}_{\text{GCL}}$ (see Appendix A.3 for more details). Moreover, tracking the evolution of uncertainty values offers an interpretable perspective on how each task contributes to the overall pre-training process, enabling a deeper understanding of the model’s learning dynamics and task interactions.
3.4. Combination with Downstream Feature Encoder
Application on Text-attributed Graphs. Text-attributed graphs consist of nodes associated with textual descriptions and edges representing their relations. Language models are commonly employed to capture both textual and structural information. In this setting, GFSE can be seamlessly integrated to inject structural awareness into textual representations. Given the generated $\mathbf{P}^{L}\in\mathbb{R}^{N\times d_{e}}$, a lightweight MLP is trained to project $\mathbf{P}^{L}$ into the embedding space of the language model. The projected PSEs are prepended as soft tokens to the textual inputs, thereby enriching them with graph topology cues. These structure-augmented tokens are then processed by downstream LLMs, improving their understanding and reasoning on graph-related tasks. The approach, involving training a lightweight MLP and fine-tuning LLM with Parameter-Efficient Fine-Tuning (PEFT) techniques such as LoRA (hu2021lora), makes it scalable and efficient for large-scale text-attributed graph applications.
3.5. Computational Complexity
GFSE’s complexity comprises two parts: pre-computation of self-supervision labels and pre-training. Computing the shortest path distance between node pairs results in the time complexity of $\mathcal{O}(|E|+|V|\log|V|)$ with node set $V$ and edge set $E$. A brute-force implementation of the subgraph isomorphism counting of fixed size $t$ is $\mathcal{O}(|V|^{t})$. We consider the graphlets with at most 5 nodes. One can also choose special graphlet types, e.g., paths, cycles, and triangles, which can be efficiently enumerated (giscard2019general). Approximating and scalable algorithms can be further used to accelerate this pre-processing step (fu2024desco; ying2020neural; pashanasangi2020efficiently). For pre-training, the complexity is $\mathcal{O}(|V|^{2})$ for full attention computation and $\mathcal{O}(d|V||E|)$ for initial encoding computation. See runtime evaluations in Appendix B.3.
4. Experiments
We first evaluate the pre-training performance in Sec. 4.2 and empirically assess the expressiveness of GFSE on synthetic datasets in Sec. 4.3. We then evaluate GFSE in a wide range of downstream graph learning tasks in Sec. 4.4. Specifically, we conduct experiments with LLMs on text-attributed graphs in Sec. 4.5.
4.1. Pre-Training Setup
Dataset. We utilize diverse datasets for pretraining, ensuring a broad spectrum of graph structures and scales, including MolPCBA, MolHIV, MNIST, peptides, ogbn-proteins, Pokec, ogbn-arxiv and ogbn-product (wu2018moleculenet; Bhatia16; mikolov2013distributed; szklarczyk2019string; chiang2019cluster; takac2012data; dwivedi2023benchmarking; dwivedi2022LRGB). These datasets cover several real-world graph domains, such as social networks, academic networks, etc. We refer to Table 8 in Appendix A.1 for detailed dataset statistics. For large-scale graphs, we first partition them into sets of subgraphs by the METIS algorithm (karypis1997metis) to handle scalability issues. Training samples from different datasets are mixed and randomly shuffled to form a large-scale pre-training dataset.
Pre-training Setting. The pre-training stage is conducted on the standard train/validation/test splits of the pre-training datasets. The initial encoding dimension $d$ is set as 8. We adopt GIN (xu2018powerful) as the message-passing layer in the GPS and adopt 8 GPS layers with 8 heads and 128 hidden dimensions. The output dimension is 64 by default. We use Adam as the optimizer with an initial learning rate of $0.001$ and the batch size is 256. The maximum training epochs is 100. An early stopping strategy is used to mitigate overfitting. The pre-training is implemented on the NVIDIA A40 48GB GPU. We refer to Appendix A.4 for more details.
4.2. Pre-training Evaluation
GFSE is pre-trained with four self-supervised learning tasks. We iteratively change the message passing layers (e.g., GatedGCN (bresson2017residual), GCN (kipf2016semi) and GIN (xu2018powerful)) and replace the biased attention with traditional self-attention in the default GFSE architecture. We evaluate the pre-training performance with different architectures as shown in Figure 3. Accuracy is used to measure community detection and graph contrastive learning tasks, indicating the proportion of node (graph) pairs that are correctly predicted. MSE and MAE are used for shortest path distance and motif counting tasks. See more details in Appendix A.2. We observe a consistent performance boost when combining biased attention with GIN, which serves as the default configuration of GFSE.
Figure 3. Pre-training performance with different architectures. TF stands for transformer.
4.3. Expressiveness Evaluation
We empirically evaluate the structure-awareness of the positional and structural encoding (PSE) generated by GFSE on Triangle (knyazev2019understanding), Pattern, and Cluster (dwivedi2023benchmarking) , three datasets that require discerning intricate graph topologies. We evaluate the performance boost brought by GFSE in comparison to traditional RWSE and LapPE. We test on various downstream graph learning models, including MLP, GIN (xu2018powerful), transformer (vaswani2017attention) and GPS (rampavsek2022recipe). We further compare with the learning-based PSE approach: GPSE (liu2023graph). As shown in Table 2, GFSE generates expressive and robust PSE that consistently improves the base model’s performance, whereas other structural encodings exhibit considerable variation across different datasets or different base models. Notably, the performance boost brought by GFSE is particularly significant for the Transformer on the Triangle dataset. This demonstrates that GFSE has a stronger enhancement effect on models that originally lack structural bias.
Table 2. Test accuracy ($\%$) enhanced by different positional and structural encoding. The results are averaged over five random seeds. The best results in each dataset are bolded.
@@ 4.3. Expressiveness Evaluation
4.4. Downstream Evaluation
4.4.1. Evaluation on Pre-training Domains
We conduct a comprehensive evaluation of GFSE on real-world graph datasets. We adhere to the experimental setting and hyper-parameters established by previous works (rampavsek2022recipe) to implement base models. We augment the initial node features with GFSE and evaluate the performance in downstream tasks, comparing it against established structural encoding methods including RWSE, LapPE, and GPSE (liu2023graph). Evaluation results on citation networks (e.g., ogbn-Arxiv, PubMed), and image graphs (e.g., MNIST and CIFAR10) are given in Table 3, where PubMed and CIFAR10 are out of the pre-training dataset. More results are provided in Table 10 in Appendix B.2.
Table 3. Test Accuracy $(\%)$ on ogbn-Arxiv, PubMed, MNIST and CIFAR10.
ogbn-arxiv PubMed MNIST CIFAR10
GateGCN $71.69_{\pm 0.21}$ $76.86_{\pm 0.41}$ $97.34_{\pm 0.14}$ $67.31_{\pm 0.31}$
TF+LapPE $5.86_{\pm 0.00}$ $66.27_{\pm 0.46}$ $96.95_{\pm 0.38}$ $69.01_{\pm 0.61}$
TF+RWSE $5.86_{\pm 0.00}$ $64.43_{\pm 0.37}$ $97.81_{\pm 0.58}$ $70.70_{\pm 0.45}$
TF+GPSE $21.56_{\pm 2.74}$ $65.89_{\pm 0.14}$ $97.78_{\pm 0.32}$ $69.57_{\pm 0.16}$
GPS $70.68_{\pm 0.71}$ $74.26_{\pm 0.60}$ $98.05_{\pm 0.12}$ $71.49_{\pm 0.35}$
GPS+GFSE $\mathbf{72.30}_{\pm 0.13}$ $\mathbf{75.13}_{\pm 0.35}$ $98.15_{\pm 0.46}$ $\mathbf{74.11}_{\pm 0.93}$
Results. From Table 3, we observe that the integration of GFSE achieves state-of-the-art performance in 10 out of 12 evaluated cases. This enhancement is not confined to a specific domain; it is evident across diverse tasks, including node classification on large-scale citation networks and graph-based image classification. Notably, GFSE consistently outperforms both traditional hand-crafted encodings, which show inconsistent benefits, and the existing pre-trained baseline, GPSE. The dramatic performance leap observed when applying GFSE to the standard Transformer is particularly telling, as it highlights our model’s ability to inject essential structural inductive biases into architecturally general models, enabling them to effectively reason over complex graph data.
4.4.2. Evaluation on Molecular Graphs
We use the molecular datasets, namely Tox21, Sider, BBBP, and MUV (hu2020open) as a downstream task. We evaluate the effectiveness of GFSE under two settings: (1) In the training-from-scratch setting, we directly concatenate GFSE’s PSE with the raw node features to create new input features. This augmented representation is then fed into a randomly initialized model from the beginning of training. We take GINE (xu2018powerful) and GPS (rampavsek2022recipe) as our backbone. (2) In the fine-tuning setting, we assess GFSE’s ability to enhance pre-trained models by concatenating the node encodings obtained from a pre-trained model with the GFSE’s PSE. The concatenated features are then fed into the final read-out layers for prediction. During fine-tuning, the parameters of the pre-trained model and read-out layers are updated. We select GraphMAE (hou2022graphmae) and MoleBERT (xia2022mole) as the pre-trained backbones and compare with other baselines without structural encoding, namely SSP (hu2019strategies), GraphLoG (xu2021self), GraphCL (you2020graph). Refer to Appendix B.1 for more implementation details.